# Unmasking Deepfakes: Leveraging Augmentations and Features Variability for Deepfake Speech Detection

### v1

@@ Abstract
The detection of deepfake speech has become increasingly challenging with the rapid evolution of deepfake technologies. In this paper, we propose a hybrid architecture for deepfake speech detection, combining a self-supervised learning framework for feature extraction with a classifier head to form an end-to-end model. Our approach incorporates both audio-level and feature-level augmentation techniques. Specifically, we introduce and analyze various masking strategies for augmenting raw audio spectrograms and for enhancing feature representations during training. We incorporate compression augmentations during the pretraining phase of the feature extractor to address the limitations of small, single-language datasets. We evaluate the model on the ASVSpoof5 (ASVSpoof 2024) challenge, achieving state-of-the-art results in Track 1 under closed conditions with an Equal Error Rate of 4.37%. By employing different pretrained feature extractors, the model achieves an enhanced EER of 3.39%. Our model demonstrates robust performance against unseen deepfake attacks and exhibits strong generalization across different codecs.
@@
journal: Speech Communication
@@ 1.1 Deepfake Speech Detection
Deepfake speech poses serious security concerns across various fields, including cybersecurity, law enforcement, and military operations. Synthetic audio can be exploited for misinformation, impersonation, and fraud, necessitating robust detection techniques to ensure the integrity of audio-based communications. Initial detection methods centered on traditional feature extraction techniques, such as Mel-frequency cepstral coefficients (MFCC) and spectrogram analysis, which capture essential acoustic properties to help differentiate genuine from synthetic speech [1, 2, 3, 4]. These traditional features were often combined with classifiers like support vector machines (SVMs) and Gaussian mixture models (GMMs), achieving initial successes in synthetic speech detection through the modeling of fundamental spectral patterns [5, 6]. Deep learning approaches, including convolutional neural networks (CNNs) and recurrent neural networks (RNNs), have advanced deepfake speech detection. CNNs, particularly ResNet models, capture local acoustic features [4, 7], while RNNs like LSTMs model temporal dependencies for improved classification [8]. Hybrid CNN-RNN architectures combine spatial and sequential features effectively [9].
The ASVSpoof challenge series [10, 11] has become a critical benchmark for developing and evaluating anti-spoofing techniques in Automatic Speaker Verification (ASV) systems, adapting over time to address increasingly sophisticated deepfake threats. The latest iteration, ASVSpoof5, consists of two tracks: Track 1 focuses on standalone deepfake speech detection, independent of ASV systems, while Track 2 integrates spoof detection within ASV pipelines to mirror real-world applications. ASVSpoof5 includes diverse and challenging datasets with various synthetic methods, codecs, and speaker variations, evaluated by Equal Error Rate (EER) and additional metrics to measure detection accuracy. By continuously raising the standard for spoof detection, ASVSpoof encourages advancements in robust, generalizable solutions to counter deepfakes, which are critical in high-stakes fields such as cybersecurity and law enforcement [12].
@@ 1.2 Speech Features
Audio features play a critical role in deepfake speech detection by providing a more structured representation of the raw audio signal, which is inherently a one-dimensional vector. Extracting meaningful features from audio, such as spectral, temporal, or statistical properties, enables models to better capture patterns and anomalies associated with deepfake generation processes
Spectrograms are a classic feature extraction technique that provide a time-frequency representation of the audio signal. By displaying the amplitude of frequencies over time, spectrograms enable models to capture not only the phonetic content but also fine-grained details of the audio, including artifacts introduced by deepfake generation methods such as speech synthesis and voice conversion [13]. Spectrograms have proven effective in deepfake detection, as mentioned in [14] and [15], CNN-based architectures applied to spectrogram inputs can significantly differentiate between genuine and spoofed speech. However, despite their success, the use of spectrograms may be approaching a performance ceiling. Recent advances suggest that further improvements in deepfake detection may require more sophisticated feature extraction methods, as the complexity of spoofing attacks continues to increase.
Wav2Vec, a deep learning-based feature extractor introduced in [16], was originally designed for automatic speech recognition (ASR) tasks. Wav2Vec 2.0, a refined version [17], leverages self-supervised learning to generate high-level speech representations directly from raw audio. Although initially developed for ASR, recent research has demonstrated the efficacy of Wav2Vec features in deepfake detection tasks [18, 19]. By utilizing self-supervised learning, Wav2Vec captures rich contextual information from speech, which is essential for identifying subtle anomalies in deepfake speech that traditional methods may miss. These representations, being task-agnostic and capable of generalizing across diverse conditions, have shown promise in enhancing deepfake detection systems against both known and novel attacks.
@@ 1.3 Speech Augmentations
Data augmentation techniques play a pivotal role in enhancing the robustness of deepfake speech detection models. By introducing variations to the training data, augmentation methods aim to simulate real-world conditions, enabling the models to learn generalized patterns that can better handle unseen, adversarial, or out-of-distribution inputs. Common speech augmentation techniques include pitch shifting, time-stretching, noise addition, and speed perturbation. These methods help models adapt to different speaker characteristics, environmental conditions, and recording qualities [3, 20].
@@ 1.4 Main Contributions
This work presents a robust model for deepfake speech detection, trained end-to-end on a limited dataset from the ASVSpoof5 dataset, achieving SoTA results in Track 1 under the closed conditions. We specifically focus on data augmentation and training techniques that enable the development of a resilient model despite the limitations of the dataset. The key contributions of this paper are as follows:
@@
    We introduce novel augmentation strategies applied at multiple stages of the audio processing pipeline. We propose the use of raw audio augmentations, which are first applied to the raw audio signal, converted to a spectrogram, and then processed as input to the model. Furthermore, we extend this approach by introducing feature-level augmentations that are applied directly to the output of the feature extractor, in the middle of the pipeline.
@@
    We integrate audio compression augmentations into the pretraining phase of the feature extractor to strengthen its robustness. This approach is particularly important in the context of self-supervised learning, especially when working with small, single-language datasets. By exposing the model to compression artifacts during pretraining, we introduce variability into the input data, which is critical for self-supervised learning. This enhances the model’s ability to generalize and improves its adaptability to new, unseen conditions.
@@
    We propose a hybrid architecture that combines Wav2Vec 2.0 structure as feature extraction with a ResNet structure as classification head. This architecture is trained using the power of self-supervised learning for feature extraction pertaining and supervised learning for training the model end-to-end for deepfake speech detection.
@@ 2 Method
This section outlines the methodology we employed, beginning with our proposed hybrid architecture, which combines a trainable feature extraction module (Wav2Vec2) with a ResNet34 classification head. Our approach incorporates augmentation strategies applied at different stages of the pipeline to enhance model performance. For raw audio, we employ MaskedSpec, low-pass filtering (LPF), and compression augmentation. Compression augmentations, in particular, are utilized both during the pretraining phase of the feature extractor and in the end-to-end training process. At the feature level, our work introduces MaskedFeature and feature normalization.
Figure 1: Illustration of our method: raw audio serves as the input, followed by raw-audio augmentations. The augmented audio is processed through a feature extractor, after which the extracted features undergo additional augmentations. Finally, the classification head processes the enhanced features to generate predictions.
2.1 Hybrid model
We propose a hybrid model that combines a trainable feature extractor with a classification head, designed to achieve joint optimization of feature extraction and classification. The training process comprises two phases: (1) a self-supervised pretraining phase for the feature extractor, enabling it to learn robust and transferable latent representations from raw audio data, and (2) an end-to-end supervised training phase for the entire hybrid model, ensuring that the extracted features align with task-specific classification objectives. As illustrated in Figure 1, augmentations are applied at multiple stages during the end-to-end training process, targeting both raw audio inputs and extracted features. The hybrid design integrates two sub-models, with each sub-model benefiting from augmentation strategies.
Leveraging the success of Wav2Vec 2.0 in speech recognition tasks, we adopted it as the feature extractor in our architecture. This choice is motivated by Wav2Vec 2.0’s ability to capture meaningful audio representations without the need for extensive labeled data, thanks to its self-supervised learning paradigm. For the classification head, we employed ResNet34, a deep convolutional neural network renowned for its robust performance in image classification and its proven adaptability for audio classification tasks, including deepfake detection using traditional frequency-domain features.
@@ 2.2 Augmentation Strategies
To address the challenges posed by unseen spoof attacks and varying codecs, we suggest augmentation strategies that aim to introduce variability during training, enabling the model to generalize better and adapt to diverse conditions  [^1] https://github.com/InbalRim/MaskedSpec. Our methodology incorporates augmentations at two distinct stages of the processing pipeline to address different aspects of corruption. First, augmentations are applied directly to the raw audio signals, modifying the dataset itself. By introducing such corruptions at the raw data level, the model is exposed to a broader range of input variations, improving its resilience. Second, augmentations are applied to the classifier input, focusing on the audio features and intermediate representations derived from the raw data. These feature-level augmentations simulate distortions that occur in processed audio, helping the model learn to recognize and handle nuanced variations in audio representations.
@@ 2.2.1 Raw Audio Augmentations
Figure 2: Visualization of different mask types used in the experiments: Squares, Bands, Singles, and Gauss.
@@
where $x(n)$ is the input audio, $w(n-m)$ represents the window function, $m$ and $k$ are the time and frequency bins, and $N$ is the frame size.
@@
Low-pass filtering (LPF) is a fundamental signal processing technique often employed to simulate the loss of high-frequency information in various audio applications. In the context of deepfake speech detection, LPF is particularly effective in modeling the compression or recording artifacts that frequently occur in real-world scenarios. By attenuating the higher-frequency components of the audio signal, LPF mimics the effects of common distortions, such as those introduced by lossy audio codecs or imperfect recording conditions. This augmentation method plays a crucial role in improving the robustness and generalization ability of the model. Additionally, the randomization of cutoff frequencies used in LPF further diversifies the training data, exposing the model to a broader range of possible high-frequency losses. This, in turn, enhances the model’s ability to generalize across different audio conditions. We perform LPF using 1D convolution operation applied to the audio signal $x[n]$ with a filter kernel $h[n]$, which acts as the low-pass filter. The filtered output signal $y[n]$ is given by the following convolution formula:
$y[n]=\sum_{m=0}^{M-1}x[n-m]h[m]$ (4)
The filter kernel, $h[m]$, is constructed using a sine function modulated by the cutoff frequency, $f_{\text{cutoff}}$, and smoothed with a Hamming window, $\text{window}[m]$, to minimize spectral leakage. The formula for $h[m]$ is given by:
$h[m]=\frac{\sin(2\pi f_{\text{cutoff}}m)}{\pi m}\cdot\text{window}[m],$
where a special case is applied at $m=0$ to prevent division by zero:
$h[0]=2\cdot f_{\text{cutoff}}.$
The use of the Hamming window reduces edge discontinuities, ensuring a smooth and stable frequency response. This design effectively attenuates frequencies above the cutoff, preserving the desired range and improving overall filter performance.
In order to address the challenges posed by real-world audio distortions, particularly from commonly used lossy compression, we incorporated a compression-decompression (codec) augmentation into our pipeline. Lossy compression algorithms such as MP3 and M4A are widely used for audio storage and streaming, due to their ability to reduce file sizes by discarding some of the less perceptible audio information. However, the compression process often introduces artifacts such as quantization noise, distortion, and loss of high-frequency content, which can negatively impact audio quality. To simulate the effects of these distortions, we applied the augmentation by encoding and decoding the raw audio at various bitrates. This encoding process generates artifacts unique to the compression algorithm used, such as the introduction of audible noise and spectral smearing, which occur due to the limited bitrate and reduced signal fidelity. In our experiments, we used the compression augmentation in both the end-to-end supervised training step and the pretraining self-supervised phase. In the end-to-end training, the model is exposed to audio samples subjected to compression and decompression at varying bitrates, compelling it to learn to detect deepfake speech even when compression artifacts are present. For self-supervised pretraining, training relies on a diverse and extensive dataset to ensure that the model can learn meaningful representations from a broad range of input variations. The variability introduced by compression artifacts further enriches the dataset, providing additional diversity and improving the model’s ability to generalize. By exposing the model to a wider variety of noise and distortion types, the pretraining process is strengthened, leading to better performance when the model is later fine-tuned for specific tasks, such as deepfake detection.
@@ 2.2.2 Feature Augmentations
To further expose the model to potential feature artifacts and improve its generalization ability, we employed an augmentation strategy similar to MaskedSpec, but applied directly to the latent representations produced by the feature extractor. This augmentation is referred to as MaskedFeature. The primary goal of this technique is to force the model to rely on broader, more generalized patterns in the data, rather than focusing on specific, localized details that might be overly sensitive to noise or distortions. In MaskedFeature, portions of the feature space are occluded during training, which simulates the presence of distortions or information loss in the latent representations. By masking these parts of the feature space, the model learns to adapt to a more holistic view of the input data, promoting robustness and resilience. This approach is particularly beneficial in scenarios where certain features of the input may be corrupted or unavailable, such as in deepfake detection tasks where some parts of the signal may be intentionally manipulated. We also investigate the use of various mask shapes for MaskedFeature, aiming to further explore how different patterns of occlusion influence the model’s ability to generalize.
Feature normalization was implemented to scale the extracted feature values into a consistent range, such as $[-1,1]$. Normalization minimizes the dominance of individual features, promoting balanced learning and reducing sensitivity to recording inconsistencies or device-specific biases. This step is critical for ensuring robustness across varied input conditions.

### v2

@@ Abstract
Deepfake speech detection presents a growing challenge as generative audio technologies continue to advance. We propose a hybrid training framework that advances detection performance through novel augmentation strategies. First, we introduce a dual-stage masking approach that operates both at the spectrogram level (MaskedSpec) and within the latent feature space (MaskedFeature), providing complementary regularization that improves tolerance to localized distortions and enhances generalization learning. Second, we introduce compression-aware strategy during self-supervised to increase variability in low-resource scenarios while preserving the integrity of learned representations, thereby improving the suitability of pretrained features for deepfake detection. The framework integrates a learnable self-supervised feature extractor with a ResNet classification head in a unified training pipeline, enabling joint adaptation of acoustic representations and discriminative patterns. On the ASVspoof5 Challenge (Track 1), the system achieves state-of-the-art results with an Equal Error Rate (EER) of 4.08% under closed conditions, further reduced to 2.71% through fusion of models with diverse pretrained feature extractors. when trained on ASVspoof2019, our system obtaining leading performance on the ASVspoof2019 evaluation set (0.18% EER) and the ASVspoof2021 DF task (2.92% EER).
@@
journal: Computer Speech & Language
@@ 1.1 Deepfake Speech Detection
Deepfake speech poses serious security concerns across various fields, including cybersecurity, law enforcement, and military operations. Synthetic audio can be exploited for misinformation, impersonation, and fraud, necessitating detection techniques to ensure the integrity of audio-based communications. Early detection systems relied on handcrafted features such as Mel-frequency cepstral coefficients (MFCCs) and spectrograms, typically processed by classifiers like support vector machines (SVMs) or Gaussian mixture models (GMMs) [1, 2, 3, 4]. These approaches offered moderate success but lacked robustness against evolving generative deepfake techniques. More recent methods have adopted convolutional neural networks (CNNs), which can effectively capture local time-frequency patterns and subtle artifacts left by generative models. Architectures such as ResNet [4, 5] and SENet [6] have shown strong performance on benchmark datasets, benefiting from their ability to model fine-grained acoustic distortions.
The ASVSpoof challenge series [7, 8] has become a critical benchmark for evaluating deepfake detection systems. The latest edition, ASVSpoof5 (ASVSpoof2024), includes two tracks: Track 1 focuses on standalone deepfake speech detection, independent of ASV systems, while Track 2 integrates detection within ASV pipelines [9]. ASVSpoof5 introduces diverse spoofing techniques, codecs, and speaker variations to reflect practical deployment scenarios. Performance is typically evaluated using metrics such as Equal Error Rate (EER), encouraging the development of detection systems that generalize across unseen attack types.
@@ 1.2 Speech Features
Feature extraction remains critical to effective deepfake detection. Traditional representations such as spectrograms offer a time-frequency view of the signal [10] and are widely used as CNN inputs. While these features have been instrumental in early systems [11, 12], the use of spectrograms may be approaching a performance ceiling. Recent advances suggest that further improvements in deepfake detection may require more sophisticated feature extraction methods, as the complexity of spoofing attacks continues to increase.
Recent studies have explored the use of pre-trained models for speech representation learning, particularly Wav2Vec 2.0 [13] and WavLM [14]. These models were originally developed for tasks such as automatic speech recognition (ASR) and speaker verification, and are trained on large-scale unlabeled speech corpora using masked prediction objectives. While not explicitly designed for deepfake detection, their learned representations have been shown to be effective in this context, as they capture rich acoustic and speaker-discriminative features that can help identify subtle artifacts introduced by synthetic speech generation [6, 15].
@@ 1.3 Speech Augmentations
Data augmentation techniques play a pivotal role in stability and broad applicability of deepfake speech detection models. By introducing variations to the training data, augmentation methods aim to simulate real-world conditions, enabling the models to learn generalized patterns that can better handle unseen, adversarial, or out-of-distribution inputs. Common speech augmentation techniques include pitch shifting, time-stretching, noise addition, and speed perturbation. These methods help models adapt to different speaker characteristics, environmental conditions, and recording qualities [3, 16].
@@
Another general-purpose augmentation technique is RawBoost [6], which applies perturbations directly in the waveform domain, such as amplitude scaling, reverberation, and filtering. Additionally, low-pass filtering (LPF) has been shown to enhance detection performance by attenuating high-frequency components, which often contain residual artifacts introduced during speech synthesis.
@@ 1.4 Main Contributions
This work presents a deepfake speech detection model trained end-to-end for the ASVSpoof5 challenge, achieving state-of-the-art performance in Track 1 under the closed condition setting. The study focuses on augmentation and training strategies designed to enhance the model’s ability to perform well despite the limited size and linguistic scope of the training data. The primary contributions of this work are summarized below:
@@
    Two-stage masking framework: We design an augmentation strategy that combines spectral masking (MaskedSpec) and latent-space masking (MaskedFeature). Multiple mask types are evaluated, and the best-performing configurations consistently improve robustness and provide additional gains when combined with established methods such as RawBoost and compression augmentation.
@@
    Compression-aware self-supervised pretraining: Introduces lossy codec perturbations directly into the self-supervised pretraining stage, unlike prior work that applies them only during supervised fine-tuning. This strategy strengthens the resilience of self-supervised representation learning for deepfake detection, particularly when training data is limited, by increasing data diversity without disrupting the underlying audio structure.
@@
    Hybrid training framework: We present a novel integration of a self-supervised, learnable feature extractor with a convolutional classifier (ResNet34). While each component is well established individually, our framework couples them within a unified hybrid training process and optimizes them jointly in an end-to-end scheme. This design enables task-specific adaptation of both low-level acoustic representations and high-level discriminative features.
@@ 2 Method
This section outlines the methodology we employed, beginning with our proposed hybrid pipeline, which combines a trainable feature extraction module (Wav2Vec2) with a classification head (ResNet34). Our approach incorporates augmentation strategies applied at different stages of the pipeline to enhance model performance. For raw audio, we employ MaskedSpec, low-pass filtering (LPF), rawboost and compression augmentation. Compression augmentations, in particular, are utilized both during the pretraining phase of the feature extractor and in the end-to-end training process. At the feature level, our work introduces MaskedFeature and feature normalization.
Figure 1: Overview of the proposed hybrid training framework. In the first stage (top, dark-shaded blocks), raw audio is used to pretrain the feature extractor using a self-supervised objective. In the second stage (bottom, light-shaded blocks), the model is trained end-to-end for the task of deepfake speech detection. Augmentations are applied at multiple points in the pipeline to boost tolerance to variability and to enrich the training signal under low-resource condition. The feature extractor is initialized with weights from the pretraining phase, as indicated by the downward arrow.
2.1 Hybrid Training Pipeline
We propose a hybrid training framework for deepfake speech detection that integrates self-supervised pretraining with supervised end-to-end training. As illustrated in Figure 1, the system consists of a trainable feature extractor and a classification head, with augmentations applied during both stages. In the first phase, the feature extractor is pretrained on unlabeled audio using a self-supervised objective. To enhance robustness in low-resource conditions, we introduce compression-based augmentations at this stage, promoting invariance in learned representations and improving feature transferability across domains. In the supervised stage, the pretrained feature extractor is initialized with the learned weights and jointly fine-tuned with the classification head using labeled data. This stage incorporates additional raw audio augmentations, including low-pass filtering, RawBoost perturbations, and spectrogram masking, as well as feature-level augmentations such as normalization and feature masking. These augmentations improve generalization by exposing the model to signal variations encountered in real-world audio. We use Wav2Vec 2.0 as the feature extractor and ResNet34 as the classification head, leveraging their respective strengths in acoustic representation and deep pattern modeling. The end-to-end training strategy enables the model to adapt both low-level and high-level representations to the task of detecting synthetic speech. An overview of the full pipeline is shown in Figure 1.
@@ 2.2 Augmentation Strategies
To enhance generalization to unseen deepfake techniques and codec conditions, we employ a comprehensive augmentation strategy applied at different stages of training  [^1] https://github.com/InbalRim/MaskedSpec.
During supervised training, we introduce a two-stage masking framework that operates at both the waveform and feature levels. These masking strategies are conceptually grounded in dropout theory, where masking acts as structured regularization. At the spectral level, masks occlude frequency regions, forcing the model to reconstruct missing information from context. At the feature level, masking suppresses localized latent activations, encouraging distributed and noise-tolerant embeddings.
In the self-supervised stage, we propose compression-aware learning by injecting lossy codec artifacts directly into the input. This approach incorporates codec variability directly into the representation learning phase, serving as a natural form of data augmentation that expands the effective training set without introducing artificial distortions that might mislead the feature extractor. In low-resource, single-language scenarios, this strategy provides a principled way to improve resilience for deepfake detection while preserving the integrity of the underlying speech signal.
Beyond these core contributions, we also evaluate a range of additional augmentations to assess their individual and combined effects. Together, these methods form a layered augmentation framework that promotes robustness to diverse perturbations and distributional shifts.
@@ 2.2.1 Raw Audio Augmentations
Figure 2: Visualization of the different mask types used in our experiments: Squares, Bands, Singles, and Gauss. White regions indicate the masking value ($\mu_{\text{stft}}$), while black regions represent the original, unmasked values.
@@
where $x(n)$ is the input audio, w(nm)w(n-m) represents the window function, $m$ and $k$ are the time and frequency bins, and $N$ is the frame size.
@@
Low-pass filtering (LPF) was used as a lightweight augmentation to simulate high-frequency loss commonly introduced by codecs or suboptimal recording conditions. By randomly varying the cutoff frequency, this technique injects spectral variability, helping the model generalize to audio degraded in different ways. LPF was implemented via 1D convolution with a Hamming-windowed sinc kernel:
$h[m]=\frac{\sin(2\pi f_{\text{cutoff}}m)}{\pi m}\cdot\text{window}[m],\quad h[0]=2\cdot f_{\text{cutoff}}.$ (4)
To increase the diversity of distortion types during training, we incorporate RawBoost as an additional waveform-level augmentation. RawBoost introduces a sequence of audio perturbations, including amplitude scaling, filtering, and reverberation, to simulate realistic acoustic degradations commonly found in practical scenarios [6]. We adopt the configuration recommended by the original authors, using the prescribed combination of distortion functions.
To improve robustness against real-world audio distortions, we incorporate an additional compression-decompression (codec) augmentation into both the self-supervised pretraining and supervised end-to-end training phases. Lossy codecs such as MP3 and M4A, commonly used for storage and streaming, introduce artifacts including quantization noise, spectral smearing, and high-frequency attenuation. These distortions can hinder detection performance, especially in mismatched conditions. Our augmentation simulates such effects by encoding and decoding audio at varying bitrates, enriching the dataset with realistic degradation. In the supervised stage, this exposes the model to codec-induced variability, helping it learn discriminative features that are robust to compression artifacts. During self-supervised pretraining, this additional variation improves representation learning by broadening the acoustic diversity encountered by the feature extractor. To maintain computational feasibility under ASVSpoof5 challenge constraints, we selected a limited yet representative set of compression formats and bitrates. This design choice was guided by prior work [5], which demonstrated that training with a small but diverse set of compressions can generalize well to unseen codec conditions without incurring prohibitive computational cost.
@@ 2.2.2 Feature Augmentations
To further expose the model to potential feature artifacts and improve its generalization ability, we employed an augmentation strategy similar to MaskedSpec, but applied directly to the latent representations produced by the feature extractor. This augmentation is referred to as MaskedFeature. The primary goal of this technique is to force the model to rely on broader, more generalized patterns in the data, rather than focusing on specific, localized details that might be overly sensitive to noise or distortions. In MaskedFeature, portions of the feature space are occluded during training, which simulates the presence of distortions or information loss in the latent representations. By masking these parts of the feature space, the model learns to adapt to a more holistic view of the input data, promoting robustness and resilience. This approach is beneficial in scenarios where certain features of the input may be corrupted or unavailable, such as in deepfake detection tasks where some parts of the signal may be intentionally manipulated. We also investigate the use of various mask shapes for MaskedFeature, aiming to further explore how different patterns of occlusion influence the model’s ability to generalize.
Feature normalization was implemented to scale the extracted feature values into a consistent range, such as [1,1][-1,1]. Normalization minimizes the dominance of individual features, promoting balanced learning and reducing sensitivity to recording inconsistencies or device-specific biases. This step is critical for ensuring robustness across varied input conditions.